# Adaptive Intelligence: leveraging insights from adaptive behavior in animals to build flexible AI systems

### v2

@@
I predict we are just on the cusp of an influx of new advances in (neuro)science due to advances in AI. For example, large-language models (LLMs) are already starting to appear in behavioral neuroscience (71) for digitizing and quantifying actions in video-based data. New vision-language, hierarchical, and multi-task foundation models are surely coming. For neural analysis, already several groups are actively working on building more unified encoders that can not only be used for brain-machine interfaces, but they themselves could serve as models of the brain (70, 29, 72, 73, 15). This is sure to give rise to new approaches to extract meaningful computational principles from neural dynamics, and these new learning rules have great potential to directly influence how we train, optimize, and deploy machine learning systems.
@@
Training and learning in artificial systems
Today, most AI applications in production revolve around a train-test-deploy cycle, and therefore are inherently not adaptive. One curates a dataset, which can be lab-project scale, or in the case of generative pretrained transformers (GPTs) nearly all the open-source content of the Internet, then split this into a train and test fraction. Most efforts benefit from leaving out-of-distribution (OOD) data [^1] OOD is defined as data that is significantly different than the training data. in the test set for a realistic measure of how well the trained model generalizes. If generalization is not needed – perhaps you are training a specific model for gait analysis in a specific laboratory setting – then the train/test split is often a random subset of the original dataset. But for many applications, training “once” and deploying a robust model is ideal. So how do you build a robust, generalizable model that can handle changes in the world? While the remarkable progress of GPTs in LLMs and the discovery of scaling laws has massively accelerated progress in AI, we aim for something better – more adaptive. What has been tried? This is where, at minimum, the sub-fields of continual learning (lifelong learning) and in-context learning come into view.
Continual learning aims to address the challenge of developing models that can learn incrementally from a continuous stream of data without forgetting previously acquired knowledge. Traditional machine learning models often face catastrophic forgetting (74) when retrained on new data, but new advancements have introduced new techniques like elastic weight consolidation (75, 76), synaptic intelligence, and memory-replay (77, 78) to mitigate these issues.
While not directly brain-inspired, Elastic Weight Consolidation (EWC) works by constraining certain parameters to reduce interference between tasks, effectively mitigating catastrophic forgetting (75). More recent work as been adapting EWC to fine-tuned self-supervised models, which has yielded performance improvements on tasks that have biased datasets by maintaining knowledge of previous tasks better than traditional methods (76, 79). If we attempt to map this to neural dynamics, this might be related to the specialization of circuits. Namely, the “parameters” of certain areas could be fixed (such as primary receptors), while others could be more plastic (such as hippocampus and neocortex).
On the other hand, synaptic intelligence is a learning rule deeply inspired by brain plasticity (80). The authors developed intelligent synapses that adaptively store information, minimizing the forgetting of previously learned tasks while acquiring new ones. This approach mimics biological neural networks that balance plasticity and stability, enabling continual learning in artificial neural networks. Notably, this method can perform as well as EWC but can be performed online. Now with modern scaleable compute and architectures, this type of learning rule could be pivotal for building more adaptive systems.
@@
Another brain-inspired machine learning algorithm has been the development of memory-replay (81, 82), which mimics the memory systems in the hippocampus (Figure 3a).  [^2] Notably, memory-replay also, perhaps indirectly, build on the memory-inspired system of the Hopfield Network, which is a type recurrent neural network designed for content-addressable memory, resembling a spin glass system (83), whose invention also won John Hopfield a Nobel Prize in 2024. Memory-replay aims to have a memory bank of already-learned actions (or images or tokens, etc.) that can be “replayed” (interweaved) into the training data in order to limit catastrophic forgetting (Figure 3b). This became a popular technique in reinforcement learning, and more recently, in computer vision and in LLMs. This training approach can also be used during active learning when new data (unlabeled, pseudo-labeled or labeled) can be used to fine-tune a model. If the input data continues to morph out-of-distribution (OOD) then using memory-replay can be leveraged to be sure the prior performance can be achieved.
@@
Several memory-replay-based approaches in LLMs emerged after the release of ChatGPT in November 2022 (with it’s impressive language abilities yet limited content window, i.e., lack of memory), such as AmadeusGPT (71), MemGPT (84), and Voyager (85). AmadeusGPT introduced a short- and long-term memory where specific key-words could be quickly recalled in order to overcome token limits (i.e., the 4096 in GPT-3.5), while MemGPT used a vector database to construct a persistent memory to maintain context across interactions, and Voyager iteratively refined its skills by replaying feedback from past actions in games such as Minecraft.
Figure 4: Foundation Models and Adaptive Agents: paths to build models for neuroscientific inference. (a) One path to building generalizable models is to input diverse data, tokenize it, and train collectively with self-supervised learning. (b) Example of Vision-Language Models (VLMs) that interweave text and images. Inspired by, and the Cat image is from, Alayrac et al. (86). (c) A proposal is to leverage/build robust specialized encoders first, then have an LLM-agent oversee accuracy and pass forward only reliable models for joint latent-space optimization. In this way, the larger model need not be taken offline but rather be systematically updated when errors are detected in input encoders. Genetic and neural icons adapted from scidraw.io; mouse images are from the Allen Brain Observatory and Lauer et al. (87).
@@
Lastly, one of the most biologically inspired and grounded advances has been in the development of spiking neural networks (SNNs) in the 1990’s. Called the “third generation of neural networks” (88), SNNs are unique in their ability to model a time-dependent spike process, making them not only biologically more plausible, but also highly energy efficient and can be used on edge-computing devices. Notably, SNNs naturally lend themselves to hardware accelerated devices and neuromorphic computing. Platforms such as Intel’s Loihi, IBM’s TrueNorth, and SpiNNaker provide hardware specifically designed to run SNNs efficiently (89). Moreover, several early work highlights their modularity and utility for multi-area modeling (90), and how multi-area SNNs could be used with semantic pointers for multi-area weighting of information (91). However, historically they have been difficult to train and scale, but that’s rapidly changing and new methods to transfer information from ANNs to SNNs are emerging  (92).
@@
How can we take inspiration from adaptive behavior and the brain to build new approaches (93)? A core component of adaptive behavior is having already learned priors – these aforementioned internal models – that can be dynamically called upon. These priors clearly require data-at-scale to form – this “data” can be baked into neural circuits via the genome, or learned throughout life. An example of genetic priors would be the innate ability to suckle at birth across many species, or how a newborn horse to stand and walk in a few hours after birth, while learned would be a motor skill such as playing a musical instrument. Another data-source is the massive amount of unsupervised sensory and motor stimuli we rapidly accumulate across development. But how do we turn data into internal models?
@@
While many questions remain about how these models are implemented in neural dynamics, there is increasing evidence of core computations (such as prediction errors (94, 41)) across brain regions, and there are architectural constraints on these systems that stem from brain-body dynamics. Brain regions are often based on primary sensors: vision, olfaction, touch, proprioception, audition, and so forth. Yet, as Mountcastle argued, it’s less about a “sensory” or “motor” space, but rather a function of inputs and outputs, and by extension, inter-connectivity of regions. Thus, as I will argue below, we should take inspiration from building specialized nodes that have core computations in the right reference framework that are interconnected in order to act.
@@
One current trend in machine learning is for large, unified Foundation Models that generalize and can be used for transfer learning in downstream tasks. The “Foundation Model” idea for natural language processing takes large-scale unlabeled data and learns a joint representation using tokenization and self-supervised learning (95) (Figure 4a). Other approaches, such as Flamingo, generalize this for multi-modal data. Flamingo is a Visual Language Model (VLM) that integrates pretrained vision and language models, processing mixed visual-text data and handling images or videos as inputs (86) (Figure 4b). This could be extended for more modalities, as recently done for robotics (96).
@@
While I believe this is exciting path for models that show better generalization, I believe we can push further to build smartly adaptive systems. We should take inspiration from our brain’s specialization and build series of smartly interconnected specialist models that can then be adapted. Namely, just as all sensory/motor information gets converted to spikes – a shared computing space and is an efficient digitization of analog signals – notably, this information is computed upon within specific brain regions. Just as a cortical column in primary visual cortex processes a particular region of the visual field and has specific coding properties related to visual features (like texture or edges), and the cortical circuit in the sensorimotor region computes in egocentric 3D kinematic, force, or even muscle space (38, 36).
@@
Moreover, across the animal kingdom we see specialization in brains everywhere: from the remarkably high-resolution of hawk vision, to bat echolocation, to the twelve cones of the mantis shrimp visual system. They aren’t sitting a single tangled mess of weights and nodes, but rather cleverly interconnected neural circuits that have specific tuning to collectively solve many tasks. While perhaps eventually a single model could reach artificial superhuman intelligence (ASI), we are far from this reality.
Thus, I argue we should also focus on building agents that are imbued with adaptive algorithms found in animal intelligence, and not only build unified monolithic Foundation Models. There is emerging work on artificial agents that I believe will make this possible. Agents that act as operation systems are one promising avenue. They can be a series of LLMs (which is currently being explored (84, 71)), but also a series of VLMs  (97), or jointly optimized encoders that individually are first pretrained (think mini-Foundation Models, such as an encoder for animal behavior (78) or unified neural region-specific pretrained encoders (70)). Then, these series of models could leverage the ‘learning’ techniques discussed above, such as memory-replay or synaptic intelligence. This is also highly likely to drive innovations, as it is also not clear that having many, independent losses is optimal.
@@
How could this work? Imagine we have a series of core-expert models (aka nodes): one for segmenting video, one for extracting poses, another for ingesting text plus images (VLMs), and one for extracting dynamics from neural spikes (Figure 4c). Each could be first experts in their specific domain, then linked via a shared latent space (even in a common spike-like framework) and optimized together. This could have a human-level language interaction agent (such as a LLM) where at inference-time the multi-modal input data could also be queried with natural language, but importantly, monitor the sub-modules (much like a prefrontal cortex). Then, we could dynamically “lock” models that show robustness on out-of-distribution samples and robustness to adversarial attacks. These nodes don’t then need to be trained (adapted) until they are deemed no longer robust. This proposal would be akin to the cortex–basal ganglia loop that has shown to be instrumental in learning and habit formation (98). Namely, these nodes could be in states of “skill learning” or “habituation” (frozen) and be constantly monitored for robustness (Figure 4c). When robust levels fall, then continual learning, memory-replay, pseudo-labeling or injection of new high-quality labeled data could be added without taking the entire system offline. There has been a surge of exciting work on building OOD detection modules (99, 100) and explainable attribution methods in time-series and images (101, 102, 103, 104, 105) that could make this an attractive path forward.
@@
Another approach, closely linked to robustness detection, could be specifically adding in prediction errors into the LLMs. Effectively, prediction errors give a signed teaching signal when an OOD sample is detected. The development of predictive-coding inspired networks already show promise (106, 107, 108), and now with modern transformer architectures or hybrid CNN-transformers, perhaps we can leverage prediction errors in order to do not only in-context learning, but signal which nodes need a new model – akin to how the brain must decide to online change a motor command vs. update an internal model. Here, we can take inspiration from the motor system (109). A one-time error means “in-context learn” and change your ongoing motor command [^3] While neurally it remains to be seen if such rapid learning (on the order of ms) can cause synaptic changes, it is doubtful that even behaviorally-time-dependent-plasticity would be fast enough., but a repetitive error means update your internal model (Figure 1c). If the incoming perturbation cannot be solved with online learning or model adaptation, a new internal model would then need to be created (which takes more time and can be completed “offline”). The neural schema in prefrontal cortex discovered by El-Gaby et al. (20) could inspire new approaches for adaptive memory-replay and new model building as well (Figure 1b).
@@
Lastly, taking the cellular diversity seriously in neural networks might drive innovations. SNNs have both excitatory and inhibitory, and even neuro-modulatory units. Can we merge the power of large transformers and the scale of data animals receive, with the cleverness of cortical circuit and/or midbrain (Figure 2)? Perhaps gating mechanisms akin to transformer blocks (Figure 4b) plus with signed (excitatory or inhibitory) tokenization and masking (perhaps mimicking biological time-delays) can lead to new innovations in adaptive artificial systems.
@@
As we strive to move beyond traditional AI towards building truly adaptive intelligence, the key lies in integration of insights from biological systems. Inherent to these ideas is that these adaptive agent systems can be put into robotic hardware. These embodied agents can then be systematically tested in true sensory-motor systems. Understanding how animals naturally adapt and rapidly learn in changing environments can guide the development of embodied AI agents that are more flexible. Taking deeper inspiration from nature in the form of adaptive behavior, specialization of species, and canonical computations in the brain, I believe, will be critical to build adaptive intelligence. By leveraging the latest findings in systems neuroscience and drawing inspiration from specialized neural computations, I hope we can design adaptive algorithms that truly mirror the sophistication of natural intelligence.

### v3

@@ Introduction
I predict we are just on the cusp of an influx of new advances in (neuro)science due to advances in AI. For example, large-language models (LLMs) are already starting to appear in behavioral neuroscience for digitizing and quantifying actions in video-based data (86). New vision-language, hierarchical, and multi-task foundation models are surely coming (87). For neural analysis, already several groups are actively working on building more unified encoders that can not only be used for brain-machine interfaces, but they themselves could serve as models of the brain (85, 39, 88, 89, 18, 90). This is sure to give rise to new approaches to extract meaningful computational principles from neural dynamics, and these new learning rules have great potential to directly influence how we train, optimize, and deploy machine learning systems.
Training and learning in artificial systems
For understanding how neuroscience is poised to influence modern AI , here I provide some background on modern AI approaches. Today, most AI applications in production revolve around a train-test-deploy cycle, and therefore are inherently not adaptive. A dataset is curated, which can be lab-project scale (or in the case of generative pretrained transformers (GPTs), this would include nearly all the open-source content of the Internet), and then split into a train-and-test fraction. Most efforts benefit from leaving out-of-distribution (OOD) data in the test set for a realistic measure of how well the trained model generalizes. OOD is defined as data that is significantly different than the training data. If generalization is not needed – perhaps you are training a specific model for gait analysis in a specific laboratory setting – then the train/test split is often a random subset of the original dataset. But for many applications, training once and deploying a robust model is ideal.
How do we build a robust, generalizable model that can handle changes in the real-world? While the remarkable progress of GPTs in LLMs and the discovery of scaling laws has massively accelerated progress in AI, we aim for something better – more adaptive. What has been tried? This is where, at minimum, the sub-fields of continual learning (lifelong learning) and in-context learning come into view.
Current approaches to adapting models
Continual learning is the task of having a neural network learn a new series of tasks of the same modality, such as computer vision tasks, over time. It aims to address the challenge of developing models that can learn incrementally from a continuous stream of data without forgetting previously acquired knowledge. This is something that biological brains excel at - we typically don’t forget older information as we learn new skills. Yet, traditional machine learning models often face catastrophic forgetting (91) when retrained on new data, but new advances in continual learning have introduced techniques such as local module composition (92), knowledge distillation (93), elastic weight consolidation (EWC) (94, 95), synaptic intelligence, and memory-replay (96, 97) to mitigate these issues. For a more extensive discussions I point the readers to 98, 99.
Although not directly brain-inspired, Elastic Weight Consolidation (EWC) works by constraining certain parameters to reduce interference between tasks, effectively mitigating catastrophic forgetting (94). More recent work has adapted EWC to fine-tune self-supervised models, which has yielded performance improvements on tasks that have biased datasets by maintaining knowledge of previous tasks better than traditional methods (95, 100). If we attempt to map this to neural dynamics, this might be related to the specialization of circuits. Namely, the “parameters” of certain areas could be fixed (such as primary receptors), whereas others could be more plastic (such as hippocampus and neocortex).
On the other hand, synaptic intelligence is a learning rule deeply inspired by brain plasticity (101). The authors developed intelligent synapses that adaptively store information, minimizing the forgetting of previously learned tasks while acquiring new ones. This approach mimics biological neural networks that balance plasticity and stability, enabling continual learning in artificial neural networks. Notably, this method can perform as well as EWC but can be performed online. Now, with modern scalable computing and architectures, this type of learning rule could be pivotal for building more adaptive systems.
@@
Another brain-inspired machine learning algorithm for learning has been the development of memory-replay (102, 103), which mimics the memory systems in the hippocampus (Figure 3a). Notably, memory-replay also, perhaps indirectly, is built on the memory-inspired system of the Hopfield Network, which is a type recurrent neural network designed for content-addressable memory, resembling a spin glass system (104), whose invention also won John Hopfield a Nobel Prize in 2024. Memory-replay aims to have a memory bank of already learned actions (or images or tokens, etc.) that can be “replayed” (interweaved) into the training data in order to limit catastrophic forgetting (Figure 3b). This became a popular technique in reinforcement learning, and more recently, in computer vision and in LLMs. This training approach can also be used during active learning when new data (unlabeled, pseudo-labeled or labeled) can be used to fine-tune a model. If the input data continues to morph out-of-distribution (OOD) then using memory-replay can be leveraged to be sure the prior performance can be achieved.
@@
Several memory-replay-based approaches in LLMs emerged after the release of ChatGPT in November 2022 (with it’s impressive language abilities yet limited content window, i.e., lack of memory), such as AmadeusGPT (86), MemGPT (105), and Voyager (106). AmadeusGPT introduced a short- and long-term memory where specific keywords could be quickly recalled in order to overcome token limits (i.e., the 4096 in GPT-3.5), while MemGPT used a vector database to construct a persistent memory to maintain context across interactions, and Voyager iteratively refined its skills by replaying feedback from past actions in games such as Minecraft.
Learning with spiking neural networks
@@
Lastly, one of the most biologically inspired and grounded advances has been in the development of spiking neural networks (SNNs) in the 1990’s. Called the “third generation of neural networks” (107), SNNs are unique in their ability to model a time-dependent spike process, making them not only biologically more plausible, but also highly energy efficient and can be used on edge-computing devices. Notably, SNNs naturally lend themselves to hardware accelerated devices and neuromorphic computing. Platforms such as Intel’s Loihi, IBM’s TrueNorth, and SpiNNaker provide hardware specifically designed to run SNNs efficiently (108). Moreover, several early work highlights their modularity and utility for multi-area modeling (109), and how multi-area SNNs could be used with semantic pointers for multi-area weighting of information (110). However, historically they have been difficult to train and scale, but that is rapidly changing and new methods to transfer information from ANNs to SNNs are emerging  (111) and/or to directly compute the required gradients from SNNs (112).
@@
How can we take inspiration from adaptive behavior and the brain to build new AI approaches (11)? A core component of adaptive behavior is having already learned priors – these aforementioned internal models – that can be dynamically called upon. These priors clearly require data-at-scale to form – this “data” can be baked into neural circuits via the genome, or learned throughout life. An example of genetic priors would be the innate ability to suckle at birth across many species, or how a newborn horse stands and walks in a few hours after birth, while a non-genetic learned prior would be a motor skill such as playing a musical instrument. Another data source is the massive amount of unsupervised sensory and motor stimuli we rapidly accumulate across development and life. But how do we turn data into internal models?
@@
While many questions remain about how these models are implemented in neural dynamics, there is increasing evidence of core computations such as the previously discussed prediction errors (113, 51) across brain regions, which shape and update internal models. Can we use these prediction errors as teaching signals in neural networks? Moreover, these prediction errors are anatomically defined, imposing architectural constraints on these systems that stem from brain-body dynamics. Can we leverage this feature in adaptive AI systems? Importantly, these prediction errors are canonical and act in specific sensory reference frameworks (such as vision or audition). As Mountcastle noted, it is less about “sensory” or “motor" spaces but rather a function of the data input/output, and by extension the shared architecture and computations. Thus, as I will argue below, we should take inspiration from building specialized nodes that have core computations in the right reference framework that use prediction errors to smartly update modules, much like the brain uses prediction error to update internal models.
Foundation Models
First, let us consider what is currently the top approach for building generalizable models whose aim is often to ingest several data modalities (such as text and video) and output either. The trend in machine learning is to build unified “Foundation Models” that generalize to unseen data or can be used for transfer learning in downstream tasks (114). The Foundation Model idea for natural language processing takes large-scale unlabeled data and learns a joint representation within a model (a transformer) using tokenization and self-supervised learning (114) (Figure 4a). The most infamous of this type is ChatGPT and other GPT models (such as the Gemini model series (115) or DeepSeek (116)). In neuroscience, several Foundation Models are also emerging for specific domains, such as for animal behavior analysis (97) and for modeling the visual systems of mice (90).
@@
Other emerging and prominent approaches are generalizing this approach to multi-modal data (117, 118). One such example is Flamingo, a Visual Language Model (VLM) that integrates pretrained vision and language models, processing mixed visual-text data and handling images or videos as inputs (117) (Figure 4b). This could be extended to more modalities, as recently done for robotics (119). This is a rapidly advancing area of research that will surely influence neuroscience in the coming years.
@@
While I believe this is an exciting path for models that show better generalization, I believe we can push further to build smartly adaptive, agentic systems. We should take inspiration from our brain’s specialization and build series of smartly interconnected specialist models that can then be adapted. Namely, just as all sensory/motor information gets converted to spikes – a shared computing space this is an efficient digitization of analog signals – notably, this information is computed upon within specific brain regions. Just as a cortical column in the primary visual cortex processes a particular region of the visual field and has specific coding properties related to visual features (like texture or edges), and the cortical circuit in the sensorimotor region computes in egocentric 3D kinematic, force, or even muscle space (48, 46).
@@
Moreover, across the animal kingdom we see specialization in brains everywhere: from the remarkably high resolution of hawk vision, to bat echolocation, to the twelve cones of the mantis shrimp visual system. The neurons that are not sitting in a single tangled mess of weights and nodes (akin to a single Foundation Model), but rather cleverly interconnected into neural circuits that have specific tuning to collectively solve many tasks. Perhaps we should take inspiration from these specialized systems versus aiming for a single model that could reach artificial superhuman intelligence. Thus, we should also focus on building smart, adaptive agentic systems of specialized models.
Figure 4: Foundation Models and Adaptive Agents. (a) One path to building generalizable models is to input diverse data, tokenize it, and train collectively with self-supervised learning. (b) Example of Vision-Language Models (VLMs) that interweave text and images. Inspired by, and the Cat image is from, Alayrac et al. (117). (c) My proposal is to leverage robust generalist foundation-model encoders that a domain specific information who’s outputs can be tokenized. For example, “pose estimation" can output skeletons of animals, “genomic" can be the DNA (or protein) sequence, “neural" could be tokenized spikes, and “vision-language" can be video captioning and actions (i..e, a higher level abstraction of behavior). In this example the assumption is there is a link between the genome, the animal phenotype, and the neural code. Critically though to my proposal is to then have an LLM-based agentic system oversee accuracy per encoder and pass forward only reliable outputs for joint latent-space optimization and joint decoder. In this way, the larger model need not be taken offline but rather be systematically updated when errors are detected in various encoders outputs. (d) Building on Bommasani et al. (114), agentic systems can leverage Foundation Models and advances in ML and deep learning in order to move beyond the emergence of functionality and into adaptive intelligent systems. Genetic and neural icons adapted from scidraw.io; mouse images are from the Allen Institute and Lauer et al. (120).
AI Agents
There is emerging work that I believe will make building AI agents (agentic systems) that are imbued with adaptive algorithms found in biologically intelligent systems possible. Agents that act as operating systems are one promising avenue. They can be a series of LLMs (which is currently being explored (105, 86)), but also a series of VLMs and other multi-modal language models (121), which work together to ingest data, select the appropriate task-specific encoder, and link these together with downstream processing scripts. Concretely, an example of this is video behavioral analysis, where an agentic system must select the best pose estimation model for the animal type in the video, perform semantic segmentation with another model, and then perform task-programming to create ethograms (86).
To develop more adaptive agentic systems, domain-specific encoder models are essential to achieve optimal performance. As in the example above, I propose an agentic system whose architecture relies on a series of encoder modules, yet adds two new aspects (Figure 4c). One, the agentic system that relies on a series of encoder modules, i.e., one for extracting poses, another for ingesting text plus images (VLMs), and one for extracting dynamics from neural spikes, is mediated by prediction errors from a decoder. Two, the outputs of each specialist encoder is jointly optimized in a latent space, which can then both learn joint representations that may be inherently more brain-like, but also a more powerful than a single Foundation Model without specialist nodes. This approach is domain-agnostic but draws on neuroscience to enhance adaptability.
@@
How could this work? Encoders would be individually pretrained (think domain specific Foundation Models, such as an encoder for animal behavior (97) or unified neural region-specific pretrained encoders (85)) and their outputs are jointly optimized into another transformer (Figure 4c). The issue becomes that while each encoder model starts as an expert, over time they need to learn more information, and the question is (1) how to sense that they need to undergo learning, and (2) how to adapt them. I propose that we could dynamically “lock” encoder models that show robustness on out-of-distribution samples and robustness to adversarial attacks; implying that they continually need monitored, by a dedicated LLM node, to provide trustworthy outputs. Therefore these encoders don’t need to be trained (adapted) until they are deemed no longer robust. This process itself I see as similar to the cortex-basal ganglia loop that is instrumental in learning and habit formation (122). Namely, these encoders could be in states of “skill learning” or “habituation” (frozen) and be constantly monitored for robustness (Figure 4c). When robustness levels fall below an acceptable threshold, then continual learning, memory-replay, pseudo-labeling or injection of new high-quality labeled data could be added without taking the entire agentic system offline, just the encoder(s) that is not trustworthy. There has been a surge of exciting work on building OOD detection modules (123, 124) and explainable attribution methods in time-series and images (125, 126, 127, 128, 129) that could make this an attractive path forward.
@@
We can also use a neuroscience-inspired approach by specifically adding prediction errors into the LLM decoder that are monitoring each encoder. Effectively, prediction errors provide a signed teaching signal when an uncertain prediction is detected. The development of predictive-coding inspired networks already show promise (130, 131, 132), and now with modern transformer architectures we could leverage prediction errors in order to do not only then do in-context learning within the encoder online, but signal which encoders need unlocked and updated – akin to how the brain must decide to online change a motor command vs. update an internal model. Here, I take inspiration from the motor system (133). A one-time error means “in-context learn” and change your ongoing motor command, but a repetitive error means update your internal model (Figure 1c). If the incoming perturbation cannot be solved with online adaptation, a new internal model would then need to be created, which takes more time and can be completed “offline” in the unlocked model. Then methods such as continual learning, pseudo-labeling and memory-replay, or synaptic intelligence could be used to adapt the model online. Moreover, new discoveries in neuroscience should inspire new online learning techniques. The neural schema in prefrontal cortex discovered by El-Gaby et al. (13) could inspire new approaches for adaptive memory-replay and new model building as well (Figure 1b). Models taken offline would use standard techniques such as gradient decent to fine-tune the model on the data that produced a prediction error (and therefore a drop in robustness).
The second aspect of my proposal is to jointly optimize a downstream joint-latent space that takes as inputs the output tokens from each specialist encoder. This could provide an embedding that has a richer representation of each sub-task vs. jointly optimizing on raw input data as in Foundation Models. Importantly, this jointly optimized space can have it’s own LLM agent, with it’s own prediction errors, that also provides human natural language interactions such that at inference-time the model outputs human-interpretable reasoning, aiding in transparency, trustworthiness, and interpretability.
@@
Lastly, independent of the agentic system, taking the cellular diversity found in the brain seriously in neural networks should drive innovations in architecture design. SNNs have both excitatory and inhibitory, and even neuro-modulatory units compared to current transformers. Can we merge the power of large transformers and the scale of data animals receive, with the diverse cell types that produce neural computations (Figure 2)? Perhaps gating mechanisms akin to transformer blocks (Figure 4b) that add a signed (excitatory or inhibitory) tokenization and masking with biological time-delays can lead to new innovations in adaptive artificial systems.
@@
As we strive to move beyond traditional AI towards building truly adaptive intelligence (Figure 4d), a key lies in the integration of insights from biological systems. I argue that adaptability will emerge from incorporating neuroscience-inspired principles, particularly internal models that leverage prediction error-based updating that refines internal representations based on discrepancies between expected and observed inputs, and from the brain-like modularity that structures the system with functionally distinct yet interconnected encoders, akin to sensory and motor modules that differ in data types but share an architecture. By integrating these mechanisms, agentic systems could achieve greater robustness and generalization across tasks, making them more effective in dynamic and uncertain environments. Inherent to these ideas is that these adaptive agentic systems can also be embodied into robotic hardware to systematically test, in fully observable sensorimotor system, their robustness and generalization. Ultimately, such advances therefore not only push progress in AI, but impact neuroscience by providing a new framework and testbed for theories of the mind.